Tickets for a football game sold for $8 and $12. If 343 tickets were sold for total receipts of $3576, how many $8 tickets and $12 tickets were sold?
Answer by mananth(16946)

$8 ticket x
$12 ticket y

1x+1y=343 .............1
8x+12y= 3576 .............2

multiply (1)by -12
Multiply (2) by 1
-12x-12 y=-4116
8x+12y= 3576
Add the two equations
-4x= -540
/-4
x=135 $8 tickets
y= 208 $12 tickets